OpenStudy (anonymous):

function rule for x: 0,1,2,3, y: 9,28,47,66

OpenStudy (zpupster):

if x is 0 y = 9 so y = x + 9 if x= 1 y = 28 28 = a(1) + 9 28 - 9 = a a= 19 so y = 19x + 9 test for x = 2 and 3